OpenStudy (anonymous):

4(-x +4)=12 multi-step equations

jigglypuff314 (jigglypuff314):

understand that the derivative property is like this: a(b+c) = a*b + a*c so when 4(-x+4) then...

OpenStudy (anonymous):

so it will be -4x+4=12 @jigglypuff314

jigglypuff314 (jigglypuff314):

4(-x+4) would be (4)*(-x) + (4)*(4)

OpenStudy (anonymous):

i dont get it @jigglypuff314

jigglypuff314 (jigglypuff314):

distributing is confusing for a lot of people... I like to think of it as what distributing means, like if I were to distribute bread I would break it into pieces and give it out so like if "a" was bread and "b" and "c" were who I gave it out to a(b + c), distributive property would be I give "a" bread to "b" and "a" bread to "c" so a*b + a*c